OpenStudy (anonymous):

what is the solution of the following system ? {-3x-2y=-12 { 9x+6y=-9 A(2,1) B no solution C (-2,-1) D infinitely many solutions

OpenStudy (dls):

-3x-2y=-12 9x+6y=-9 Multipy 1st equation by 3 -9x-6y=-24 Add them 9x+6y=-9 -9x-6y=-24 ---------- 0+0=-33 ----------- Hence no solution :)

OpenStudy (anonymous):

okay i get it now thank you :)